Look at this Table: This table was created with LYX 1.1.6fix1 on Windows 2000. No changes

where made in the sources concerning any LYX code creation!

Up to LYX 1.1.5fix2 I was always able to use LYX files on both Linux and Win32. Now a table can’t be seen as a table under Linux when written under Windows although the same LYX release is used!

Operatoren </cell></row><row topline="true" bottomline="false"

newpage="false"><cell multicolumn="0" alignment="center" valignment="top" topline="true" bottomline="false" leftline="true"

rightline="false" rotate="false" usebox="none" width="" special="">++ --

Vorzeichen ~ ! (type) . new</cell><cell multicolumn="0" alignment="center" valignment="top" topline="true" bottomline="false" leftline="true" rightline="true" rotate="false" usebox="none" width="" special="">unär (nur auf ein Argument bezogen!)</cell></row><row

topline="true" bottomline="false" newpage="false"><cell multicolumn="0" alignment="center" valignment="top" topline="true" bottomline="false" leftline="true" rightline="false" rotate="false" usebox="none" width=""

special="">* / %</cell><cell multicolumn="0" alignment="center" valignment="top" topline="true" bottomline="false" leftline="true"

rightline="true" rotate="false" usebox="none" width="" special="">von links nach rechts</cell></row><row topline="true" bottomline="false"

newpage="false"><cell multicolumn="0" alignment="center" valignment="top" topline="true" bottomline="false" leftline="true" rightline="false" rotate="false" usebox="none" width="" special="">+ -</cell><cell multicolumn="0" alignment="center" valignment="top"

topline="true" bottomline="false" leftline="true" rightline="true" rotate="false" usebox="none" width="" special="">von links nach

rechts</cell></row><row topline="true" bottomline="false" newpage="false"><cell multicolumn="0" alignment="center"

valignment="top" topline="true" bottomline="false" leftline="true" rightline="false" rotate="false" usebox="none" width="" special=""><< >> >>></cell><cell multicolumn="0" alignment="center" valignment="top"

topline="true" bottomline="false" leftline="true" rightline="true" rotate="false" usebox="none" width="" special="">von links nach

rechts</cell></row><row topline="true" bottomline="false" newpage="false"><cell multicolumn="0" alignment="center"

valignment="top" topline="true" bottomline="false" leftline="true" rightline="false" rotate="false" usebox="none" width="" special="">< <=

> >= instanceof</cell><cell multicolumn="0" alignment="center" valignment="top" topline="true" bottomline="false" leftline="true"

rightline="true" rotate="false" usebox="none" width="" special="">von links nach rechts</cell></row><row topline="true" bottomline="false"

newpage="false"><cell multicolumn="0" alignment="center" valignment="top" topline="true" bottomline="false" leftline="true"

rightline="false" rotate="false" usebox="none" width="" special="">== !=</cell><cell multicolumn="0" alignment="center" valignment="top"

topline="true" bottomline="false" leftline="true" rightline="true" rotate="false" usebox="none" width="" special="">von links nach

rechts</cell></row><row topline="true" bottomline="false" newpage="false"><cell multicolumn="0" alignment="center"

valignment="top" topline="true" bottomline="false" leftline="true"

rightline="false" rotate="false" usebox="none" width="" special="">&</cell><cell multicolumn="0" alignment="center"

valignment="top" topline="true" bottomline="false" leftline="true" rightline="true" rotate="false" usebox="none" width="" special="">von links nach rechts</cell></row><row topline="true" bottomline="false"

newpage="false"><cell multicolumn="0" alignment="center" valignment="top" topline="true" bottomline="false" leftline="true"

rightline="false" rotate="false" usebox="none" width="" special="">^</cell><cell multicolumn="0" alignment="center"

valignment="top" topline="true" bottomline="false" leftline="true" rightline="true" rotate="false" usebox="none" width="" special="">von links nach rechts</cell></row><row topline="true" bottomline="false"

newpage="false"><cell multicolumn="0" alignment="center" valignment="top" topline="true" bottomline="false" leftline="true"

rightline="false" rotate="false" usebox="none" width="" special="">|</cell><cell multicolumn="0" alignment="center"

valignment="top" topline="true" bottomline="false" leftline="true" rightline="true" rotate="false" usebox="none" width="" special="">von links nach rechts</cell></row><row topline="true" bottomline="false"

newpage="false"><cell multicolumn="0" alignment="center" valignment="top" topline="true" bottomline="false" leftline="true"

rightline="false" rotate="false" usebox="none" width="" special="">&&</cell><cell multicolumn="0" alignment="center" valignment="top" topline="true" bottomline="false" leftline="true"

rightline="true" rotate="false" usebox="none" width="" special="">von links nach rechts</cell></row><row topline="true" bottomline="false"

newpage="false"><cell multicolumn="0" alignment="center" valignment="top" topline="true" bottomline="false" leftline="true"

rightline="false" rotate="false" usebox="none" width="" special="">||</cell><cell multicolumn="0" alignment="center" valignment="top" topline="true" bottomline="false" leftline="true"

rightline="true" rotate="false" usebox="none" width="" special="">von links nach rechts</cell></row><row topline="true" bottomline="false"

newpage="false"><cell multicolumn="0" alignment="center" valignment="top" topline="true" bottomline="false" leftline="true"

rightline="false" rotate="false" usebox="none" width="" special="">?:</cell><cell multicolumn="0" alignment="center" valignment="top" topline="true" bottomline="false" leftline="true"

rightline="true" rotate="false" usebox="none" width="" special="">von rechts nach links</cell></row><row topline="true" bottomline="true"

newpage="false"><cell multicolumn="0" alignment="center" valignment="top" topline="true" bottomline="false" leftline="true"

rightline="false" rotate="false" usebox="none" width="" special="">= += -= /= %= · · · · · ·</cell><cell multicolumn="0" alignment="center" valignment="top" topline="true" bottomline="false" leftline="true"

rightline="true" rotate="false" usebox="none" width="" special="">von rechts nach links</cell></row></lyxtabular>

Tabelle 1: Priorität und Assoziativität der Operatoren 2